OpenStudy (anonymous):

Solve the equation. -1/8(x-16)1/8(x-8)=x+5

OpenStudy (anonymous):

\[\frac{ \frac{ -1 }{ 8(x-16) } }{ 8(x-8) }\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

Its 1/8(x-16) 1/8(x-8)= x+5

OpenStudy (anonymous):

\[\frac{ \frac{ -1 }{ 81(x-16) } }{8(x-8 }=x+5\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

first \[\frac{ -1 }{ (81*8)(x-16)(x-8) }=x+5\]

OpenStudy (anonymous):

\[-1=648(x+5)(x-16)(x-8)\]

OpenStudy (anonymous):

Oh I'm sorry 81 isn't in the problem. Its -1/8 and a positive 1/8.

OpenStudy (anonymous):

rewrite the equation please :)

OpenStudy (anonymous):

-1/8 (x-16) 1/8 (x-8)=x+5

OpenStudy (anonymous):

\[\frac{ -1 }{ 8(x-16)? }\times \frac{ 1 }{ 8(x-8)? }=x-5\]

OpenStudy (anonymous):

Yes but its (x+5)

OpenStudy (anonymous):

ok just replace the 648 by 64 :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then multiple the three elements and collect the similar terms

OpenStudy (anonymous):

77x+48?

OpenStudy (anonymous):

no :(